# Splitting differential equations using Galois theory

### v1

@@ 3.2. Groupoids, splitting and short exact sequences
Recall that definable maps give rise to definable morphisms of binding groups (see for example [10, Lemma 3.1]). More precisely:
Fact 3.11.
@@
Let $f:p\rightarrow f(p)$ be an $A$-definable fibration with $\mathcal{C}$-internal fibers. For any $f(a),f(b)\models f(p)$, we define the set of morphisms $\mathrm{Mor}(f(a),f(b))$ to be the set of bijections from $p_{f(a)}$ to $p_{f(b)}$ which extend to an automorphism of $\mathcal{U}$ fixing $A\cup\mathcal{C}$ pointwise. We define the groupoid associated to $f$, denoted $\mathrm{Gd}_{A}(p,f/\mathcal{C})$, to be the groupoid with set of objects $f(p)$ and morphisms the sets $\mathrm{Mor}(f(a),f(b))$. This groupoid acts naturally on $p$ in the following way: by construction, any element of $\mathrm{Mor}(f(a),f(b))$ is a bijection from $p_{f(a)}$ to $p_{f(b)}$, and we define its (partial) action on $p$ to be that bijection. The following is [7, Theorem 1.3]:
@@
The groupoid $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is isomorphic to an $A$-type-definable groupoid, and its natural action on $p$ is $A$-definable.
@@
For a proof, going from retractability to splitting is given by[7, Proposition 4.3], and the other direction is given in the proof of [8, Theorem 3.3.5]. We do not include a proof, but give a brief sketch.
Assuming $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is retractable, we have a family of map $r(f(a),f(b))$. We define an equivalence relation $E$ on $p$ as $E(x,y)$ if and only if $r(f(x),f(y))(x)=y$. The compatibility conditions of retractability exactly mean that this is an equivalence relation. The map $\pi$ is the quotient map of $E$. The other direction is similar to the right to left direction of the next lemma.
When $f(p)$ is fundamental, the connection between splitting of the short exact sequence and retractability is straightforward. This is proven in [7, Theorem 4.10], but we include a considerably streamlined proof:
@@
As $f(p)$ is fundamental, the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ is free and transitive. As $f(p)$ is weakly orthogonal to $\left\{\pi(p),\mathcal{C}\right\}$, so is the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))$, and therefore we must have $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))=\mathrm{Aut}_{A}(f(p)/\mathcal{C})$, and by Lemma 2.10 we also obtain $\mathrm{Aut}_{A}(\pi(p)/\mathcal{C},f(p))=\mathrm{Aut}_{A}(f(p)/\mathcal{C})$.
It is easy to see that $\mathrm{Aut}_{A}(\pi(p)/\mathcal{C},f(p))\times\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))<\operatorname{Im}(\widetilde{\pi}\times\widetilde{f})$: given $(\sigma,\tau)$ in the former group, we can extend each to $\widehat{\sigma},\widehat{\tau}\in\mathrm{Aut}_{A}(p/\mathcal{C})$, and then $(\widetilde{\pi}\times\widetilde{\pi})(\widehat{\sigma}\widehat{\tau})=(\sigma,\tau)$. Therefore the map $\widetilde{\pi}\times\widetilde{f}$ is an $A$-definable isomorphism between $\mathrm{Aut}_{A}(p/\mathcal{C})$ and $\mathrm{Aut}_{A}(\pi(p)/\mathcal{C})\times\mathrm{Aut}_{A}(f(p)/\mathcal{C})$. ∎
@@
The assumption that $f(p)$ is fundamental can always be obtained, without changing the short exact sequence, by taking a high enough Morley power of $p$ (see [6, Fact 2.4]). However, applying Lemma 3.15 would yield a splitting of a Morley power of $p$, which is not what we want. Note that the right to left direction of the proof only uses that $f(p)$ is fundamental to obtain the equality $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))=\mathrm{Aut}_{A}(f(p)/\mathcal{C})$. We could therefore replace the assumption that $f(p)$ is fundamental by this and obtain the same implication.
@@
Let $p\in S(A)$ be a $\mathcal{C}$-internal type and $f:p\rightarrow f(p)$ be an $A$-definable fibration. Suppose $f(p)$ is weakly $\mathcal{C}$-orthogonal and that the short exact sequence $1\rightarrow K\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})\xrightarrow{\widetilde{f}}\mathrm{Aut}_{A}(f(p)/\mathcal{C})\rightarrow 1$ is $A$-definably split via some $A$-definable section $s$. If the following holds:
($\bigstar$): for any $f(a)\models f(p)$ and $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$, if $\sigma(f(a))=f(a)$, then $s(\sigma)|_{p_{f(a)}}=\operatorname{id}$.
then $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is retractable (and in particular $f$ splits).
Before starting with the proof, it is worth pointing out what assumption ($\bigstar$) means. If $\sigma$ fixes $f(a)$, then $s(\sigma)$ fixes $p_{f(a)}$, the fiber above $f(a)$, as a set. This assumption says that it in fact fixes $p_{f(a)}$ pointwise.
@@
Let $f(a),f(b)\models f(p)$ and $\sigma,\tau\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ such that $\sigma(f(a))=f(b)=\tau(f(a))$. Then $s(\sigma)|_{p_{f(a)}}$ and $s(\tau)|_{p_{f(a)}}$ are in $\mathrm{Mor}(f(a),f(b))$. By assumption $(s(\sigma)|_{p_{f(a)}})^{-1}\circ s(\tau)|_{p_{f(a)}}=\operatorname{id}|_{p_{f(a)}}$, so $s(\sigma)|_{p_{f(a)}}=s(\tau)|_{p_{f(a)}}$. Therefore, we can define, for any $f(a),f(b)\models f(p)$, the map $r(f(a),f(b))=s(\sigma)_{p_{f(a)}}$, for any $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ with $\sigma(f(a))=f(b)$. This is well-defined by the previous discussion, and it is straightforward to show that it witnesses retractability of $\mathrm{Gd}_{A}(p,f/\mathcal{C})$. ∎
Note that if $f(p)$ is fundamental, then any $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ fixing some $f(a)\models f(p)$ must be the identity. Therefore in that case, we always have (normal-$\bigstar$). If $f(p)$ is not fundamental, we find the following sufficient condition for (normal-$\bigstar$):
@@
Under the same assumptions as Lemma 3.16, if for some (any) $f(a)\models f(p)$, there is no non-trivial (i.e. with image not equal to the trivial subgroup) definable morphism from any definable subgroup of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ to $\mathrm{Aut}_{Af(a)}(p_{f(a)}/\mathcal{C})$ then assumption (normal-$\bigstar$) holds. In particular $f$ splits.
@@
To go from unipotent and tori to nilpotent, we need to remark that (normal-$\bigstar$) is preserved under direct product. More precisely:
@@
Let $p\in S(A)$ be a $\mathcal{C}$-internal type and $f:p\rightarrow f(p)$ be an $A$-definable fibration. Suppose $f(p)$ is weakly $\mathcal{C}$-orthogonal and that the short exact sequence $1\rightarrow K\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})\xrightarrow{\widetilde{f}}\mathrm{Aut}_{A}(f(p)/\mathcal{C})\rightarrow 1$ is $A$-definably split via some $A$-definable section $s$. Suppose that $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ is $A$-definably isomorphic to a direct product $H_{1}\times H_{2}$, that for all $f(a)\models f(p)$ we have $H_{1}f(a)\cap H_{2}f(a)=\{f(a)\}$, and that the induced actions of $H_{1}$ and $H_{2}$ on $f(p)$ both satisfy (normal-$\bigstar$). Then the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ on $f(p)$ satisfies (normal-$\bigstar$).
@@
We can write any $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ as $\sigma_{1}\sigma_{2}$ with $\sigma_{i}\in H_{i}$. Suppose that $\sigma(f(a))=f(a)$ for some $f(a)\models f(p)$. Then $\sigma_{1}(f(a))=\sigma_{2}^{-1}(f(a))$, therefore $\sigma_{1}(f(a))=\sigma_{2}(f(a))=f(a)$. By our (normal-$\bigstar$) assumptions, we obtain that $s(\sigma_{i})|_{p_{f(a)}}=\operatorname{id}$, and therefore $s(\sigma)|_{p_{f(a)}}=\operatorname{id}$. ∎

### v3

@@ 3.2. Groupoids, splitting and short exact sequences
Recall that definable maps give rise to definable morphisms of binding groups (see for example [12, Lemma 3.1]). More precisely:
Fact 3.8.
@@
Let $f:p\rightarrow f(p)$ be an $A$-definable fibration with $\mathcal{C}$-internal fibers. For any $f(a),f(b)\models f(p)$, we define the set of morphisms $\mathrm{Mor}(f(a),f(b))$ to be the set of bijections from $p_{f(a)}$ to $p_{f(b)}$ which extend to an automorphism of $\mathcal{U}$ fixing $A\cup\mathcal{C}$ pointwise. We define the groupoid associated to $f$, denoted $\mathrm{Gd}_{A}(p,f/\mathcal{C})$, to be the groupoid with set of objects $f(p)$ and morphisms the sets $\mathrm{Mor}(f(a),f(b))$. This groupoid acts naturally on $p$ in the following way: by construction, any element of $\mathrm{Mor}(f(a),f(b))$ is a bijection from $p_{f(a)}$ to $p_{f(b)}$, and we define its (partial) action on $p$ to be that bijection. The following is [9, Theorem 1.3]:
@@
The groupoid $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is isomorphic to an $A$-type-definable groupoid, and its natural action on $p$ is relatively $A$-definable.
@@
For a proof, going from retractability to splitting is given by[9, Proposition 4.3]. However as pointed out by the anonymous referee, the weak orthogonality of $f(p)$ to $\{\mathcal{C},\pi(p)\}$ is not proved there. The other direction is given in the proof of [10, Theorem 3.3.5], and is similar to the right to left direction of the next lemma. For these reasons, we do not give a full proof, but only a brief sketch of the left to right implication below.
Assuming $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is retractable, we have a family of map $r(f(a),f(b))$. We define an equivalence relation $E$ on $p$ as $E(x,y)$ if and only if $r(f(x),f(y))(x)=y$. The compatibility conditions of retractability exactly mean that this is an equivalence relation. The map $\pi$ is the quotient map of $E$. We need to show that $f(p)$ is weakly orthogonal to $\{\mathcal{C},\pi(p)\}$, or equivalently, that $\mathrm{Aut}_{A}(\mathcal{U}/\mathcal{C},\pi(p))$ acts transitively on $f(p)$. Pick any $f(a),f(b)\models f(p)$ and let $\sigma$ be any extension of $r(f(a),f(b))$ to an element of $\mathrm{Aut}_{A}(\mathcal{U}/\mathcal{C})$.
Since $E$ is $A$-definable, we know that $\sigma$ permutes the classes of $E$. Any $E$-class has a representant $a^{\prime}\models p_{f(a)}$, and we have that $E(a^{\prime},\sigma(a^{\prime}))$. Therefore $\sigma$ fixes the class of $a^{\prime}$ pointwise. As this is true for all classes, we get that $\sigma$ fixes $\pi(p)$ pointwise, and therefore restricts to an element of $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))$. This shows that $f(p)$ is weakly orthogonal to $\{\mathcal{C},\pi(p)\}$, and the second condition is easy to show (and proved in [9, Proposition 4.3]).
When $f(p)$ is fundamental, the connection between splitting of the short exact sequence and retractability is straightforward. This is proven in [9, Theorem 4.10], but we include a considerably streamlined proof:
@@
As $f(p)$ is fundamental, the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ is free and transitive. As $f(p)$ is weakly orthogonal to $\left\{\pi(p),\mathcal{C}\right\}$, so is the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))$, and therefore we must have $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))=\mathrm{Aut}_{A}(f(p)/\mathcal{C})$, and by Lemma 2.10 we also obtain $\mathrm{Aut}_{A}(\pi(p)/\mathcal{C},f(p))=\mathrm{Aut}_{A}(\pi(p)/\mathcal{C})$.
It is easy to see that $\mathrm{Aut}_{A}(\pi(p)/\mathcal{C},f(p))\times\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))<\operatorname{Im}(\widetilde{\pi}\times\widetilde{f})$: given $(\sigma,\tau)$ in the former group, we can extend each to $\widehat{\sigma},\widehat{\tau}\in\mathrm{Aut}_{A}(p/\mathcal{C})$, and then $(\widetilde{\pi}\times\widetilde{f})(\widehat{\sigma}\widehat{\tau})=(\sigma,\tau)$. Therefore the map $\widetilde{\pi}\times\widetilde{f}$ is an $A$-definable isomorphism between $\mathrm{Aut}_{A}(p/\mathcal{C})$ and $\mathrm{Aut}_{A}(\pi(p)/\mathcal{C})\times\mathrm{Aut}_{A}(f(p)/\mathcal{C})$. ∎
Note that even when $f(p)$ is not fundamental, if there is an $A$-definable section $s$, by Corollary 2.7, the group $\operatorname{Im}(s)$ is an $A$-definable normal subgroup of $\mathrm{Aut}_{A}(p/\mathcal{C})$, and this group is $A$-definably isomorphic to $\mathrm{Aut}_{A}(f(p)/\mathcal{C})\times\operatorname{Im}(s)$.
@@
The assumption that $f(p)$ is fundamental can always be obtained, without changing the short exact sequence, by taking a high enough Morley power of $p$ (see [8, Fact 2.4]). However, applying Lemma 3.12 would yield a splitting of a Morley power of $p$, which is not what we want. Note that the right to left direction of the proof only uses that $f(p)$ is fundamental to obtain the equality $\mathrm{Aut}_{A}(f(p)/\mathcal{C},\pi(p))=\mathrm{Aut}_{A}(f(p)/\mathcal{C})$. We could therefore replace the assumption that $f(p)$ is fundamental by this and obtain the same implication.
@@
Definition 3.13.
Let $p\in S(A)$ be a $\mathcal{C}$-internal type and $f:p\rightarrow f(p)$ be an $A$-definable fibration. Suppose $f(p)$ is weakly $\mathcal{C}$-orthogonal and that the short exact sequence
is $A$-definably split via some $A$-definable section $s$.
Let $H$ be a definable subgroup of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$. We say that $H$ satisfies ()($\bigstar$)( ★ ) if for any $f(a)\models f(p)$ and $\sigma\in H$, if $\sigma(f(a))=f(a)$, then $s(\sigma)|_{p_{f(a)}}=\operatorname{id}$.
Let us point out what ()($\bigstar$)( ★ ) means. If $\sigma$ fixes $f(a)$, then $s(\sigma)$ fixes $p_{f(a)}$, the fiber above $f(a)$, as a set. This assumption says that it in fact fixes $p_{f(a)}$ pointwise. We obtain:
@@
Let $p\in S(A)$ be a $\mathcal{C}$-internal type and $f:p\rightarrow f(p)$ be an $A$-definable fibration. Suppose $f(p)$ is weakly $\mathcal{C}$-orthogonal and that the short exact sequence $1\rightarrow K\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})\xrightarrow{\widetilde{f}}\mathrm{Aut}_{A}(f(p)/\mathcal{C})\rightarrow 1$ is $A$-definably split via some $A$-definable section $s$. If $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ satisfies ()($\bigstar$)( ★ ), then $\mathrm{Gd}_{A}(p,f/\mathcal{C})$ is retractable (and in particular $f$ splits).
@@
Let $f(a),f(b)\models f(p)$ and $\sigma,\tau\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ such that $\sigma(f(a))=f(b)=\tau(f(a))$. Then $s(\sigma)|_{p_{f(a)}}$ and $s(\tau)|_{p_{f(a)}}$ are in $\mathrm{Mor}(f(a),f(b))$. By assumption $(s(\sigma)|_{p_{f(a)}})^{-1}\circ s(\tau)|_{p_{f(a)}}=s(\sigma\circ\tau^{-1})|_{p_{f(a)}}=\operatorname{id}|_{p_{f(a)}}$, so $s(\sigma)|_{p_{f(a)}}=s(\tau)|_{p_{f(a)}}$. Therefore, we can define, for any $f(a),f(b)\models f(p)$, the map $r(f(a),f(b))=s(\sigma)_{p_{f(a)}}$, for any $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ with $\sigma(f(a))=f(b)$. This is well-defined by the previous discussion, and it is straightforward to show that it witnesses retractability of $\mathrm{Gd}_{A}(p,f/\mathcal{C})$. ∎
Note that if $f(p)$ is fundamental, then any $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ fixing some $f(a)\models f(p)$ must be the identity. Therefore in that case, we always have ($\bigstar$). If $f(p)$ is not fundamental, we find the following sufficient condition for ($\bigstar$):
@@
Under the same assumptions as Lemma 3.14, if for some (any) $f(a)\models f(p)$, there is no non-trivial (i.e. with image not equal to the trivial subgroup) definable morphism from any definable subgroup of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ to $\mathrm{Aut}_{Af(a)}(p_{f(a)}/\mathcal{C})$ then assumption ($\bigstar$) holds. In particular $f$ splits.
@@
To go from unipotent and tori to nilpotent, we need to remark that ($\bigstar$) is preserved under direct product. More precisely:
@@
Let $p\in S(A)$ be a $\mathcal{C}$-internal type and $f:p\rightarrow f(p)$ be an $A$-definable fibration. Suppose $f(p)$ is weakly $\mathcal{C}$-orthogonal and that the short exact sequence $1\rightarrow K\rightarrow\mathrm{Aut}_{A}(p/\mathcal{C})\xrightarrow{\widetilde{f}}\mathrm{Aut}_{A}(f(p)/\mathcal{C})\rightarrow 1$ is $A$-definably split via some $A$-definable section $s$. Suppose that $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ is $A$-definably isomorphic to a direct product $H_{1}\times H_{2}$, that for all $f(a)\models f(p)$ we have $H_{1}f(a)\cap H_{2}f(a)=\{f(a)\}$, and that the induced actions of $H_{1}$ and $H_{2}$ on $f(p)$ both satisfy ($\bigstar$). Then the action of $\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ on $f(p)$ satisfies ($\bigstar$).
@@
We can write any $\sigma\in\mathrm{Aut}_{A}(f(p)/\mathcal{C})$ as $\sigma_{1}\sigma_{2}$ with $\sigma_{i}\in H_{i}$. Suppose that $\sigma(f(a))=f(a)$ for some $f(a)\models f(p)$. Then $\sigma_{1}(f(a))=\sigma_{2}^{-1}(f(a))$, therefore $\sigma_{1}(f(a))=\sigma_{2}(f(a))=f(a)$. By our ($\bigstar$) assumptions, we obtain that $s(\sigma_{i})|_{p_{f(a)}}=\operatorname{id}$, and therefore $s(\sigma)|_{p_{f(a)}}=\operatorname{id}$. ∎